Lean Accounting

Westgate Inc. uses a lean manufacturing strategy to manufacture DVR (digital video recorder) players. The company manufactures DVR players through a single product cell. The budgeted conversion cost for the year is $729,300 for 1,870 production hours. Each unit requires 20 minutes of cell process time. During March, 950 DVR players were manufactured in the cell. The materials cost per unit is $62. The following summary transactions took place during March:

  1. Materials were purchased for March production.
  2. Conversion costs were applied to production.
  3. 950 DVR players were assembled and placed in finished goods.
  4. 900 DVR players were sold for $340 per unit.

a. Determine the budgeted cell conversion cost per hour. If required, round to the nearest dollar.
$ per hour

b. Determine the budgeted cell conversion cost per unit. If required, round to the nearest dollar.
$ per unit

c. Journalize the summary transactions (1)–(4) for March.

1.
2.
3.
4. Sale
4. Cost
Answer #1
a
Budgeted cell conversion cost per hour 390 per hour =729300/1870
b
Budgeted cell conversion cost per unit 130.00 per unit =390*(20/60)
c
Raw and in process inventory 58900 =950*62
     Accounts Payable 58900
Raw and in process inventory 123500 =950*130
       Conversion costs 123500
Finished goods inventory 182400 =950*(62+130)
       Raw and in process inventory 182400
Accounts Receivable 306000 =900*340
       Sales 306000
Cost of goods sold 172800 =900*(62+130)
        Finished goods inventory 172800